OpenStudy (anonymous):

Solve for x in the proportion 3 over 8 equals x over 12

OpenStudy (anonymous):

We have: \[\frac{ 3 }{ 8 } = \frac{ x }{ 12 }\]To solve this we want to start by cross multiplying. So we get: \[3 \times 12 = 8 \times x\]Now to solve for x we just divide both sides by 8: \[x=3 \times 12 / 8 = 36/8=18/4=9/2\]

OpenStudy (anonymous):

so x = 36? what next

OpenStudy (anonymous):

No, x is equal to 9/2. And you're done.

OpenStudy (anonymous):

ohhh i see now thank you

OpenStudy (anonymous):

so other wise x = 4.5?

OpenStudy (anonymous):

Correct.